Using the quadratic equation x^2 - 4x - 5 = 0, perform the following tasks:
Solve by factoring.
Solve by completing the square.
Solve by using the quadratic formula.

Answer by ikdeep(189)

x^2 - 4x - 5 = 0
Here you must know how we make factors i.e. splitting of the middle term...if you might not knew then contact me again and I will explain to you... we get
x^2 - (5-1)x - 5 = 0
solving of the brackets would give us
x^2 - 5x + 1x - 5 = 0
taking common from first two i.e. x^2 - 5x and last two i.e. 1x - 5 variables...we get
x(x-5) + 1(x-5) = 0
(x-5)(x+1) = 0
now either (x-5)= 0 or(x+1) = 0
therefore the values of are 5 and -1
hope this will help you
Please feel free to revert back for any further queries